# references.bib AIJIM: A Theoretical Model for Real-Time, Crowdsourced Environmental Journalism with AI

### v3

@@ III-D AIJIM Workflow
Hazard Identification: Users upload geotagged images via the NamicGreen app or other platforms.
@@
AI-Based Detection: In this study, the Vision Transformer-based DETR model demonstrates effective real-time hazard detection capabilities. AIJIM’s modular design, however, allows integration of various alternative AI models depending on specific requirements.
@@
Visual Hazard Validation: AI-detected hazards are visually verified by crowdsourced validators through direct image annotations, ensuring accuracy of identified hazards before generating reports.
@@
AI-Based Reporting and Textual Validation: Structured reports are automatically generated via GPT-4 (adaptable to alternative NLP models like Claude or Gemini). These AI-generated textual reports are subsequently validated by community and expert validators, ensuring compliance with journalistic and ethical standards.
III-E Core Components
III-E1 Automated Environmental Monitoring
In this study, AIJIM leverages a DETR-based Vision Transformer for efficient hazard detection, optimized using ONNX and TensorRT, allowing real-time processing of user-submitted images [Tiltack2025]. However, this is only one possible technological integration of AIJIM. The model’s modularity allows for alternative object detection models such as YOLO, Faster R-CNN, or Swin Transformer, ensuring adaptability to different hardware environments and accuracy requirements. This architecture was chosen due to its:
Superior ability to detect small and partially obscured hazards in diverse environments.
Lower computational cost compared to traditional object detection models such as YOLOv8 and Faster R-CNN.
Robustness across diverse lighting conditions, which is critical for real-world reporting scenarios [Carion2020, Tiltack2025].
To evaluate its performance, a pilot study on Mallorca (2024) analyzed 1,000 geotagged images uploaded via the NamicGreen platform. The DETR model achieved:
85.4% ± 1.8% detection accuracy, identifying 50 previously undocumented waste sites [Tiltack2025, Li2024].
Real-time inference capabilities, suitable for rapid environmental hazard detection [Tiltack2025].
III-F Stakeholder Interaction and Validation
AIJIM enhances detection reliability by overlaying identified hazards directly onto the images submitted by users. This allows community validators to review AI-detected findings visually and confirm their accuracy. The framework’s participatory approach is further supported by interactive maps that display user-generated environmental reports, enabling citizens to contribute to and monitor environmental issues in real-time, as demonstrated in the NamicGreen implementation (see Abbildung 9 in the Appendix). Within the NamicGreen ecosystem:
252 community validators actively participated in the verification process.
High agreement rates with expert annotators, confirming its reliability [Tiltack2025].
To incentivize participation, AIJIM integrates gamification elements, rewarding users with reputation scores based on validation contributions.
To further validate the practical impact of AI-assisted verification, 12 expert interviews were conducted with professionals from environmental journalism, AI development, and regulatory fields. These experts provided qualitative insights into the reliability and ethical implications of AI-driven environmental monitoring. Key findings from these interviews include:
AI enhances efficiency but requires transparency. Experts highlighted that journalists need interpretability tools such as Explainable AI (XAI) to ensure credibility in AI-generated content [GarciaTapiaMateo2025].
Bias remains a concern. AI systems can reinforce existing stereotypes if datasets are not properly curated. NamicGreen’s detection accuracy showed a 30% accuracy drop in rural settings, emphasizing the need for dataset expansion [Floridi2023].
Trust in AI-driven journalism depends on human oversight. Although AIJIM achieved an 89.7% agreement rate with expert validators, multiple experts stressed that hybrid AI-human collaboration is crucial to maintaining journalistic integrity [Verma2024].
Human-in-the-loop validation for AI-generated reports. To ensure journalistic accuracy and prevent the dissemination of unverified AI-generated content, all reports generated by AIJIM within the NamicGreen platform are initially stored as drafts. These drafts undergo manual review, where human validators can edit, refine, or reject AI-generated reports before they are published or forwarded to authorities. This process ensures compliance with EU AI Act 2025 regulations on AI accountability and transparency [Gstrein2024, Bolda2024].
III-G AI-Based Narrative Generation
In this implementation, AIJIM utilizes OpenAI’s GPT-4 API for automated environmental reporting, efficiently transforming hazard data into structured, journalistically relevant articles. Nevertheless, AIJIM’s modular architecture permits the integration of alternative NLP models according to specific contextual or ethical considerations.
A critical component of AIJIM’s automated reporting is its ability to accurately identify and classify objects within user-submitted images, such as various types of waste (e.g., plastic foil, rubber waste), as validated through the NamicGreen pilot study (see Abbildung 6 in the Appendix). These AI-generated insights subsequently form the basis for detailed environmental reports that outline potential hazards and offer actionable recommendations (see Abbildung 7 in the Appendix). The automation provided by AIJIM significantly reduces the latency in environmental hazard identification, enabling quicker interventions to emerging environmental threats.
Key features of AIJIM’s narrative generation include:
Automated news summarization derived from AI-validated hazard reports.
Context-aware content generation, dynamically adjusting tone and complexity based on audience preferences.
Integration with GIS visualization tools (e.g., Mapbox APIs) to enhance public accessibility and engagement [OpenAI2023, Tiltack2025].
Empirical testing demonstrated that AI-driven reporting reduced news production latency by 40%, enabling journalists and policymakers to respond faster to emerging environmental threats [Tiltack2025].
Additionally, expert interviews provided qualitative insights into the ethical concerns surrounding AI-generated journalism. A key issue identified was algorithmic bias in AI-generated content, which could lead to selective reporting. To mitigate this, experts recommended the integration of Explainable AI (XAI) techniques, such as SHAP (Shapley Additive Explanations), to provide transparency into AI-generated outputs [Floridi2023].
III-H Empirical Advantages of AIJIM
AIJIM demonstrates significant improvements over traditional environmental journalism methodologies. Empirical results from the Mallorca pilot study and expert evaluations indicate:
40% reduction in data collection and reporting time, enabled by real-time inference in this DETR-based implementation [Tiltack2025].
Higher accuracy and reliability due to direct validation via annotated images (89.7% agreement rate) [Tiltack2025].
Expert-backed validation confirming AIJIM’s potential to improve journalistic credibility and mitigate misinformation risks [Pareek2024].
III-I Performance Comparison with Traditional Methods
To highlight AIJIM’s methodological improvements, we comparatively evaluated AIJIM against traditional investigative journalism methods. Key comparison dimensions included data collection efficiency, hazard detection accuracy, reporting latency, scalability, and stakeholder engagement. Detailed quantitative performance metrics are reported in Section V, Table III.
IV Implementation in NamicGreen
@@
The Artificial Intelligence Journalism Integration Model (AIJIM) was first implemented in NamicGreen, an AI-powered platform designed for real-time environmental monitoring and automated reporting. This proof-of-concept deployment demonstrated AIJIM’s ability to process high-volume citizen-generated data while maintaining scalability, efficiency, and accuracy.
Technical Implementation: NamicGreen integrates multiple AI-driven components to optimize hazard detection, data validation, and automated news generation. The system employs advanced Computer Vision models optimized for high-speed, real-time image recognition, ensuring adaptability to various AI architectures [Tiltack2025].
Beyond object detection, NamicGreen incorporates a crowdsourced validation mechanism to enhance accuracy and mitigate AI biases. In the 2024 Mallorca pilot study, this approach yielded an 89.7% agreement rate with expert annotations, reinforcing trust in AI-driven environmental journalism [Tiltack2025]. Validators assess AI predictions using a weighted voting mechanism, where confidence scores influence validation priority. Ambiguous cases are escalated for expert review, refining AI-generated classifications, reducing false positives, and improving contextual accuracy [Tiltack2025]. The platform also integrates Explainable AI (XAI) techniques, such as SHAP, to provide transparency and interpretability of AI-generated classifications.
Scalability and Adaptability: AIJIM’s modular design allows NamicGreen to be deployed across different hardware and network environments. To ensure real-world applicability, the platform supports cloud-based processing for large-scale analyses and enables edge AI deployment for remote areas with limited connectivity. Additionally, NamicGreen’s reporting workflow is designed for flexibility, allowing integration with various NLP models. While this study utilizes GPT-4 for automated report generation, the system can adapt to alternative language models to address different linguistic or ethical considerations [OpenAI2023].
By combining real-time AI processing, scalable deployment strategies, and participatory validation, NamicGreen exemplifies how AIJIM can bridge the gap between automation and human oversight in environmental journalism.
IV-A System Architecture
As a proof-of-concept, NamicGreen demonstrates AIJIM’s ability to seamlessly integrate AI-driven workflows into real-world environmental journalism. Its scalable and modular architecture ensures adaptability across various deployment scenarios (Fig. 3).
The NamicGreen implementation is built upon a modular architecture, integrating seven core components that reflect AIJIM’s flexible and scalable design:
(1) User/NamicGreen Drone: Captures geotagged environmental data (images and GPS).
@@ IV-A System Architecture
(2) AI Analysis & Classification: Employs advanced CV models to classify pollution patterns in real-time, adaptable to various AI architectures.
(3) Geolocation & GIS Integration: Maps GPS data via Flask API for geospatial insights.
(4) Automated Report Generation: Generates initial reports from analyzed data using NLP models (GPT-4 in this implementation), adaptable to alternative models as needed.
@@
(5) Community Validation: Refines reports through crowdsourced feedback for enhanced accuracy.
@@
(6) Final Report & Data Export: Delivers validated reports to authorities, NGOs, and journalists.
(7) AI Feedback Loop: Continuously refines model performance using real-world data.
IV-B AI Image Recognition Module
Within NamicGreen, AIJIM employs advanced Computer Vision models optimized for rapid real-time inference. The modularity of AIJIM, however, supports alternative detection models depending on specific requirements and deployment scenarios. Trained and validated using diverse geotagged imagery, it ensures rapid and accurate detection, forwarding ambiguous cases for crowdsourced validation [Tiltack2025].
IV-C Crowdsourced Validation
While AI-powered hazard detection significantly improves speed and scalability, maintaining high accuracy remains a challenge—especially in complex real-world scenarios. To ensure reliability, NamicGreen incorporates a crowdsourced validation system, reinforcing AIJIM’s commitment to transparency and data integrity. Validators assess AI predictions using a weighted voting mechanism, where confidence scores influence validation priority. Ambiguous cases are escalated for expert review, refining AI-generated classifications, reducing false positives, and improving contextual accuracy [Tiltack2025].
IV-D Data Visualization and Automated Reporting
AIJIM, via NamicGreen, transforms raw environmental data into actionable insights using GIS-based visualization and automated reporting. Real-time geospatial heatmaps, powered by Mapbox APIs, pinpoint pollution hotspots with precision. These visualizations provide journalists with immediate insights, assist policymakers in prioritizing response efforts, and enable environmental agencies to dynamically track pollution trends. In this implementation, GPT-4-generated reports significantly reduced reporting latency compared to traditional methods [OpenAI2023, Tiltack2025]. AIJIM’s adaptable design also supports alternative NLP models, ensuring flexibility across different journalistic contexts.
IV-E Real-World Deployment: Mallorca Pilot Study
@@
The 2024 Mallorca pilot validated AIJIM’s real-world potential through NamicGreen, processing 1,000 citizen-submitted geotagged images to uncover 50 previously undocumented illegal waste sites [Tiltack2025]. This deployment demonstrated AIJIM’s efficacy and flexibility in realistic scenarios. Key outcomes include:
Latency Reduction: Achieved a 40% faster reporting cycle, enhancing real-time responsiveness.
Community Engagement: Engaged 252 active validators, scaling participation effectively.
Detection Accuracy: Significantly surpassed manual inspection methods, validating AIJIM’s reliability in real-time hazard detection [Tiltack2025].
@@
Future developments include scaling data collection (10,000 images), integrating multispectral satellite analysis, broadening model diversity (YOLO, Claude), and expanding multilingual capabilities for broader accessibility [Ivancsics2019, Tiltack2025]. AIJIM sets a new standard for real-time, ethical, and participatory environmental journalism.
V Results and Discussion
@@
The empirical evaluation of the Artificial Intelligence Journalism Integration Model (AIJIM) within the NamicGreen platform reveals its exceptional efficacy, scalability, and real-world relevance. The 2024 Mallorca pilot study quantitatively benchmarks AIJIM against traditional journalism, while expert insights illuminate its technical strengths and ethical challenges, underscoring its transformative potential in AI-driven environmental reporting [Tiltack2025].
V-A Performance Evaluation
AIJIM’s performance was rigorously validated in the 2024 Mallorca pilot, representing one of the first large-scale empirical studies to integrate real-time AI-based environmental reporting with crowdsourced verification. This ability to autonomously detect and verify hazards in near-instantaneous timeframes is essential for mitigating environmental crises that can escalate within hours, such as oil spills, illegal logging, or industrial pollution. Analyzing 1,000 citizen-generated geotagged images via the NamicGreen app, a 5-fold cross-validation (95% confidence interval, ) ensured robust, reproducible results. A detailed comparison of AIJIM’s implementation using a Vision Transformer-based DETR model with traditional methods is provided in Table III in Appendix A, though AIJIM’s modular framework allows for alternative models like YOLO or Faster R-CNN to be employed in different contexts.
In this implementation, the DETR model demonstrated robust precision and recall, achieving strong mAP scores and near-instantaneous inference speeds through GPU acceleration [Tiltack2025]. AIJIM’s modular design, however, remains adaptable to alternative detection models. Crowdsourced validation further improved accuracy through high expert agreement, significantly reducing reporting latency and enhancing community engagement.
@@ V-A Performance Evaluation
V-B Stakeholder Perspectives on AIJIM
To gain qualitative insights, twelve expert interviews with professionals from journalism, AI development, and regulation were conducted. These discussions highlighted AIJIM’s strengths and identified areas for further refinement (see Table IV in Appendix A).
@@ V-B Stakeholder Perspectives on AIJIM
Experts lauded AIJIM’s 85.4% precision, yet noted a 30% rural accuracy drop, echoing Section V-E1’s call for broader datasets [Tiltack2025].
@@
The 40% faster reporting was praised, but transparency concerns highlighted the need for XAI tools like SHAP and LIME to sustain credibility [Shin2021].
@@
Crowdsourcing’s 252 validators impressed, though experts urged stronger misinformation safeguards [Verma2024].
@@
AIJIM’s GDPR compliance was affirmed, yet accountability gaps under the EU AI Act 2025 persist, necessitating journalism-specific AI ethics [Gstrein2024].
V-C Addressing Expert Concerns: Future Strategies
@@
V-D Comparison with Traditional Methods
The advantages of AIJIM become even clearer when compared to conventional journalistic workflows, highlighting improvements in efficiency, accuracy, and engagement. Traditional reporting methods often require hours to days for data verification and dissemination, whereas AIJIM streamlines this process with automation and participatory validation. AIJIM’s real-time data acquisition and automation cut latency by 40%, while 252 validators quintupled engagement over passive methods (Fig. 5) [Broussard2019]. Its 85.4% precision trumps manual variability, ensuring consistent, reliable reporting [Tiltack2025, Kazaz2021]. A detailed comparison between AIJIM and traditional journalism is provided in Table V in Appendix A.
Traditional environmental journalism, while essential, remains constrained by time-intensive fieldwork, limited scalability, and high resource costs. AIJIM’s ability to autonomously detect and validate environmental hazards in real time establishes a new standard, bridging the gap between investigative rigor and technological efficiency.
V-E Challenges and Ethical Considerations
@@
V-E1 AI Bias and Generalizability
One of the key limitations observed in AIJIM’s implementation is a 30% reduction in accuracy in rural environments compared to urban areas. This discrepancy arises from multiple factors:
Limited Training Data: Rural environments often have fewer annotated datasets, leading to less model familiarity with region-specific hazard patterns [McGovern2022].
Variability in Image Conditions: Uneven lighting, dense vegetation, and limited infrastructure alter the visual characteristics of environmental hazards, making recognition harder for AIJIM [Tiltack2025].
Underrepresentation in Crowdsourcing: The lower participation of rural users in crowdsourced validation may introduce biases that reinforce incorrect classifications [Gondwe2024].
To mitigate these issues, future iterations of AIJIM will incorporate:
Expanded and balanced datasets, integrating more rural hazard samples through targeted data collection [Tiltack2025].
Adaptive learning techniques, leveraging transfer learning from satellite imagery to improve feature generalization [Li2024].
Synthetic data augmentation, generating rural-specific hazard conditions to diversify model exposure [McGovern2022].
Geo-weighted validation, where rural reports are prioritized for expert review to reduce confirmation bias in crowdsourced annotations [Gondwe2024].
These improvements will ensure AIJIM maintains high accuracy across diverse geographic regions, reducing urban-centric biases in AI-driven environmental journalism [Tiltack2025, Li2024, McGovern2022, Gondwe2024].
V-E2 Misinformation Risks
One of the main challenges in AI-generated environmental journalism is the risk of producing false positives, missed hazards, or exaggerated claims [AlZoubi2024]. These risks stem from several factors:
Bias & Generalization Issues: If the AI model is trained on unbalanced datasets, it may overemphasize certain types of hazards while neglecting others. Additionally, NLP models can sometimes generate overly confident statements based on limited data, potentially misleading readers and policymakers [Cazzamatta2025, McGovern2022, Gondwe2024].
Contextual Misinterpretation: AI models, including GPT-4, may struggle to distinguish between correlated but unrelated environmental factors, such as differentiating industrial pollution from natural algae blooms [McGovern2022].
To mitigate these misinformation risks, AIJIM incorporates the following safeguards:
Uncertainty Scoring (Planned Enhancement): Future versions of AIJIM could incorporate a confidence rating for AI-generated reports. This would allow low-confidence outputs to be flagged for manual review, reducing the risk of misinformation [Wolfe2024].
Explainable AI (XAI): The model integrates SHAP (Shapley Additive Explanations) to provide transparency into how conclusions were derived [Shin2021].
Hybrid AI-Human Verification & Fact-Checking: Before publication, AI-generated reports undergo validation by expert reviewers and crowdsourced validators. Additionally, the AI cross-references data with authoritative sources, such as environmental agencies and scientific databases, to minimize misinformation risks [Dierickx2024, Pareek2024].
These measures ensure that AI-driven journalism remains transparent, accountable, and resistant to bias, strengthening its credibility as a tool for environmental reporting [AlZoubi2024, Cazzamatta2025, McGovern2022, Gondwe2024, Wolfe2024, Dierickx2024, Shin2021, Pareek2024].
V-E3 Transparency and Accountability
To maintain journalistic integrity in AI-driven reporting, AIJIM employs Explainable AI (XAI) tools such as SHAP and LIME. These techniques provide interpretability in model decisions, ensuring that hazard classifications and automated reports remain auditable by journalists, regulators, and the public, aligning with the EU AI Act 2025 [Floridi2021].
V-E4 Fake News Prevention
AIJIM combats misinformation through automated threat detection, real-time verification, and multi-source cross-checking, reinforcing trust in AI-generated journalism [Tiltack2025].
V-E5 Regulatory and Privacy Risks
GDPR compliance is secured with AES-256 encryption and human-in-the-loop validation, though accountability gaps linger [Gstrein2024].
V-E6 Human Oversight
AIJIM balances automation with hybrid workflows, empowering journalists without replacing them [deLimaSantos2022].
V-E7 Future Governance
Federated learning, cross-cultural validation, and journalism-specific ethics will refine AIJIM’s integrity [Tiltack2025].
V-E8 Unintended Risks
Crowdsourced validation strengthens AI verification by resolving ambiguous cases (e.g., low visibility, occlusions) and mitigating dataset bias. However, challenges like misinformation and over-reliance on automation persist, necessitating continued regulatory oversight. AIJIM enhances rather than replaces journalistic integrity [AlZoubi2024].
VI Conclusion and Future Work

### v5

@@ 3.4 AIJIM Workflow
Data Collection: Users upload geotagged images via platforms like NamicGreen.
@@
AI-Based Analysis: A Vision Transformer-based DETR model enables real-time detection, with support for alternative models [34].
@@
Automated Reporting: Adaptable NLP models generate draft reports.
@@
Human Validation: Crowdsourced validators verify detections through image annotations, ensuring accuracy.
5.
Expert & Community Review: Experts and community members validate reports for ethical compliance.
6.
Publication & Distribution: Validated reports are published and distributed to stakeholders.
3.5 Core Components
3.5.1 Automated Environmental Monitoring
AIJIM employs a DETR-based Vision Transformer, optimized with ONNX and TensorRT, for real-time hazard detection [34]. Its modularity supports models like YOLO or Swin Transformer, chosen for:
Superior detection of small, obscured hazards.
Lower computational cost than YOLOv8 or Faster R-CNN.
The 2024 Mallorca pilot analyzed 1,000 images, achieving 85.4% ± 1.8% accuracy and identifying 50 waste sites [34, 24].
3.6 Stakeholder Interaction and Validation
3.7 AI-Based Narrative Generation
AIJIM’s reporting module generates narratives, reducing latency by 40% [34]. Supporting NLP models like GPT-4, Claude, Gemini, and LLaMA, it adapts to linguistic and ethical needs [29, 2, 18]. The Mallorca pilot validated object classification (e.g., plastic foil), producing actionable reports (Appendix, Abbildung 6, 7). Features include:
Dynamic summarization for tailored reports.
Context-aware tone adjustment.
GIS visualization via Mapbox APIs [34].
XAI techniques (SHAP) ensure transparency, mitigating bias, as confirmed by experts [15, 36].
3.8 Empirical Advantages of AIJIM
3.9 Performance Comparison with Traditional Methods
3.10 Explainability Architecture: CAM and LIME
AIJIM’s dual-layer explainability uses Class Activation Mapping (CAM), as illustrated in Appendix 10, for real-time heatmaps and optional LIME for detailed audits. CAM visualizes detection focus, while LIME perturbs regions for deeper insight, balancing scalability and interpretability [34].
LIME provides an additional layer of transparency by highlighting the specific regions of an image that contribute to the model’s decision. A detailed example of this process for a ”metal can” detection can be found in Appendix 11. The image in the appendix illustrates the areas of interest identified by LIME and their impact on the classification decision, offering a deeper understanding of the model’s behavior.
4 Implementation in NamicGreen
@@
The Artificial Intelligence Journalism Integration Model (AIJIM) was implemented in NamicGreen, an AI-powered platform for real-time environmental monitoring and reporting. This proof-of-concept deployment showcased AIJIM’s ability to process high-volume citizen-generated data with scalability and efficiency.
Technical Implementation: NamicGreen integrates AI-driven components for hazard detection, validation, and reporting, as detailed in Section 3. Its modular design supports cloud-based processing for large-scale analyses and edge AI deployment for remote areas with limited connectivity, ensuring real-world applicability [34]. The platform’s flexible reporting workflow adapts to various NLP models, extending beyond the GPT-4 implementation used in this study [29].
NamicGreen exemplifies AIJIM’s balance of automation and human oversight, leveraging crowdsourced validation and Explainable AI (XAI) to enhance accuracy and transparency, as validated in the 2024 Mallorca pilot (Section 4.5).
4.1 System Architecture
NamicGreen integrates seven core components, with a focus on scalable design and real-time data processing. Details on specific subsystems such as AI image recognition and crowdsourced validation can be found in Sections 3.5.1 and 3.6.
User/NamicGreen Drone: Captures geotagged images and GPS data.
@@ 4.1 System Architecture
AI Analysis & Classification: Performs real-time pollution detection, adaptable to various CV models (see Section 3.5.1).
Geolocation & GIS Integration: Maps data via Flask API for geospatial insights.
Automated Report Generation: Produces draft reports using NLP models (see Section 3.7).
@@
Community Validation: Refines outputs through crowdsourced feedback (see Section 3.6).
@@
Final Report & Data Export: Delivers validated reports to stakeholders.
AI Feedback Loop: Continuously improves model performance.
4.2 AI Image Recognition Module
4.3 Crowdsourced Validation
4.4 Data Visualization and Automated Reporting
4.5 Real-World Deployment: Mallorca Pilot Study
@@
The 2024 Mallorca pilot validated AIJIM’s potential, processing 1,000 citizen-submitted geotagged images to uncover 50 undocumented illegal waste sites. Key outcomes include:
Latency Reduction: 40% faster reporting cycle.
Community Engagement: 252 active validators.
Detection Accuracy: 85.4% ± 1.8%, surpassing manual methods [34].
@@
AIJIM’s dual explainability module (Section 3.10) deployed CAM-based overlays for real-time validation and a LIME-based API for selective expert audits, balancing transparency and scalability. Future developments include scaling to 10,000 images, integrating multispectral satellite analysis, and expanding model diversity (e.g., YOLO, Claude) and multilingual capabilities [21, 34].
5 Results and Discussion
@@
The evaluation of the Artificial Intelligence Journalism Integration Model (AIJIM) within the NamicGreen platform demonstrates its efficacy and scalability. The 2024 Mallorca pilot benchmarks AIJIM against traditional journalism, with expert insights highlighting technical and ethical considerations [34].
5.1 Performance Evaluation
The 2024 Mallorca pilot processed 1,000 citizen-generated geotagged images, detecting 50 undocumented waste sites with an accuracy of 85.4% 1.8%. Crowdsourced validation by 252 participants achieved 89.7% expert agreement, reducing reporting latency by 40% compared to traditional methods [34]. A 5-fold cross-validation (95% confidence interval, ) ensured the robustness of these results. Detailed metrics are provided in Table 3 in Appendix A.
The pilot utilized a Vision Transformer-based DETR model (Section 3.5.1) for real-time hazard detection, while crowdsourced validation (Section 3.6) significantly contributed to both model accuracy and public engagement.
@@ 5.1 Performance Evaluation
5.2 Stakeholder Perspectives
Qualitative insights from twelve expert interviews (journalism, AI, regulatory fields) are summarized in Table 4 in Appendix A.
@@ 5.2 Stakeholder Perspectives
Rapid reporting was valued, but XAI tools (SHAP, LIME) were critical for credibility [32].
@@
Crowdsourcing was effective, though misinformation safeguards were recommended [36].
@@
GDPR compliance was confirmed, but EU AI Act 2025 accountability gaps need journalism-specific ethics [19].
5.3 Addressing Expert Concerns: Future Strategies
@@
5.4 Comparison with Traditional Methods
AIJIM outperforms traditional journalism, as shown in Figure 5 and Table 5 in Appendix A. Pilot results (Section 5.1) highlight 40% faster processing and fivefold participation [4, 34, 22].
AIJIM integrates established technologies like Vision Transformers (DETR), CAM, and LIME, but its true innovation lies in the modular integration of these models into a real-time, scalable framework designed specifically for environmental journalism. By leveraging crowdsourced validation and providing explainable AI (XAI), AIJIM goes beyond the capabilities of traditional journalism models, offering greater transparency and higher engagement.
While traditional methods rely on retrospective data collection and manual verification, AIJIM utilizes real-time hazard detection, allowing for immediate reporting with improved accuracy and participation. This combination of existing AI technologies with innovative workflows marks AIJIM as a breakthrough model in the field of automated environmental journalism.
AIJIM’s real-time detection overcomes traditional journalism’s limitations by offering faster reporting, increased scalability, and greater public involvement.
5.5 Challenges and Ethical Considerations
@@
5.5.1 AI Bias and Generalizability
5.5.2 Misinformation Risks
5.5.3 Transparency and Governance
5.5.4 Unintended Risks
Crowdsourcing mitigates bias but requires oversight to prevent misinformation [1].
5.6 Explainability Outcomes
The Mallorca pilot study validated AIJIM’s effectiveness in detecting 50 undocumented waste sites from 1,000 citizen-generated geotagged images. The model achieved 85.4% accuracy, with crowdsourced validation contributing to an 89.7% expert agreement rate. This demonstrates AIJIM’s ability to process large datasets quickly, reducing reporting latency by 40% compared to traditional methods. Furthermore, the use of Vision Transformer-based DETR models ensured rapid detection, and crowdsourced validation enhanced both accuracy and engagement. These results confirm AIJIM’s potential as a scalable, real-time tool for environmental reporting.
6 Conclusion and Future Work